# Let’s Grow an Unbiased Community : Guiding the Fairness of Graphs via New Links

### v1

@@ 5.1.3. Implementation Details
[…]
@@
[…]
pokec-n $\Delta_{SP}$ () 77.3 $\pm$ 0.3 74.4 $\pm$ 2.7 77.5 $\pm$ 2.6 79.6 $\pm$ $3 \cellcolor$ lightgray79.2 $\pm$ 2.9 80.0 $\pm$ 1.4 69.2 $\pm$ 4.1
[…]
@@ 5.2. Results of Fairness Guidance
Generalization to Various GNN Architectures. In addition, to further verify the structural fairness of graphs guided by FairGuide, we evaluate the impact of fairness-guided links on two different GNN backbones, i.e. , GraphSage(hamilton2017inductive, ) and APPNP(gasteiger_predict_2019, ). Results in fairness are presented in Fig. 3. The F1 scores on different GNNs are given in Fig. 9. From Fig 3, we observe that different GNN backbones consistently achieve fairness improvements when trained on graphs guided by FairGuide. This demonstrates that FairGuide can grow fair graph structures that facilitate different GNN models.
@@ 5.5. Hyperparameter Analysis
In FairGuide, the cross-group boost rate $\beta$ and the number of community clusters C are two key hyperparameters. To investigate how the two hyperparameters affect fairness and task performance, We conduct GCN node classification task on the Pokec-n. Similar trends are also observed in other datasets. We vary $\beta$ to {0.01,0.1,1.0,4.0,10.0} and $C$ to {5,10,15,20,25}, respectively. All other settings follow the description given in Sec. 5.1. As shown in Fig. 7, the task performance F1 score remains stable across the tested hyperparameter ranges, fluctuating by only around 1–2 percentage points. For the fairness performance, FairGuide achieves a strong debiasing effect when $C\geq 10$ and $\beta\geq 1.0$. A larger number of clusters $C$ allows the model to capture more fine-grained feature information, and a large cross-group boost rate $\beta$ effectively establishes connections between different sensitive groups, which will benefit the graph structural fairness.
@@ Appendix C Algorithm of FairGuide
(a) APPNP
(b) GraphSage
Figure 9. F1 scores on different GNNs.

### v3

@@ 5.1.3. Implementation Details
Github F1 () 78.6 $\pm$ 0.2 78.0 $\pm$ 0.1 78.8 $\pm$ 0.1 78.6 $\pm$ 0.1 77.5 $\pm$ 0.2 77.5 $\pm$ 0.2 77.8 $\pm$ 0.1
AUC () 85.4 $\pm$ 0.5 84.3 $\pm$ 0.1 85.4 $\pm$ 0.1 84.8 $\pm$ 0.3 84.2 $\pm$ 0.1 84.0 $\pm$ 0.1 84.3 $\pm$ 0.1
$\Delta_{SP}$ () 12.5 $\pm$ 0.4 11.0 $\pm$ 0.2 11.9 $\pm$ 0.2 11.5 $\pm$ 0.8 11.1 $\pm$ 0.3 10.8 $\pm$ 0.3 08.6 $\pm$ 0.2
$\Delta_{EO}$ () 08.5 $\pm$ 0.5 08.3 $\pm$ 0.3 08.3 $\pm$ 0.4 8.5 $\pm$ 0.8 08.4 $\pm$ 0.3 08.3 $\pm$ 0.1 06.0 $\pm$ 0.2
pokec-n F1 () 66.8 $\pm$ 0.7 67.1 $\pm$ 0.3 66.7 $\pm$ 0.5 65.4 $\pm$ 0.4 65.4 $\pm$ 0.3 65.2 $\pm$ 0.3 66.2 $\pm$ 0.3
AUC () 75.4 $\pm$ 0.1 74.8 $\pm$ 0.1 75.4 $\pm$ 0.2 72.8 $\pm$ 0.6 74.3 $\pm$ 0.1 74.2 $\pm$ 0.2 74.9 $\pm$ 0.2
$\Delta_{SP}$ () 08.5 $\pm$ 0.7 09.2 $\pm$ 1.0 07.9 $\pm$ 0.8 03.7 $\pm$ 1.3 09.7 $\pm$ 0.4 09.6 $\pm$ 0.7 01.3 $\pm$ 0.8
$\Delta_{EO}$ () 11.9 $\pm$ 1.4 012.3 $\pm$ 1.0 11.0 $\pm$ 1.6 06.3 $\pm$ 1.5 011.7 $\pm$ 1.6 011.7 $\pm$ 0.5 03.1 $\pm$ 0.5
pokec-z F1 () 70.6 $\pm$ 0.4 70.5 $\pm$ 0.2 70.3 $\pm$ 0.6 67.1 $\pm$ 1.1 69.0 $\pm$ 0.5 69.1 $\pm$ 0.9 70.2 $\pm$ 0.4
AUC () 77.2 $\pm$ 0.1 76.8 $\pm$ 0.1 77.2 $\pm$ 0.1 75.5 $\pm$ 0.7 76.0 $\pm$ 0.1 76.0 $\pm$ 0.3 76.1 $\pm$ 0.1
$\Delta_{SP}$ () 09.1 $\pm$ 0.9 08.8 $\pm$ 0.8 09.5 $\pm$ 0.8 05.0 $\pm$ 2.0 06.4 $\pm$ 1.5 06.9 $\pm$ 2.4 03.1 $\pm$ 0.7
$\Delta_{EO}$ () 08.2 $\pm$ 1.2 08.0 $\pm$ 0.7 08.1 $\pm$ 1.1 05.1 $\pm$ 1.8 05.5 $\pm$ 2.0 07.6 $\pm$ 2.6 04.6 $\pm$ 0.4
@@
Github $\Delta_{SP}$ () 41.9 $\pm$ 1.5 35.7 $\pm$ 2.8 38.2 $\pm$ 4.0 40.2 $\pm$ 3.6 31.9 $\pm$ 2.1 36.3 $\pm$ 3.8 29.0 $\pm$ 2.2
pokec-n $\Delta_{SP}$ () 83.2 $\pm$ 3.0 76.1 $\pm$ 0.8 77.5 $\pm$ 2.6 79.6 $\pm$ 3.3 79.2 $\pm$ 2.9 80.0 $\pm$ 1.4 74.1 $\pm$ 3.4
pokec-z $\Delta_{SP}$ () 77.7 $\pm$ 0.9 79.3 $\pm$ 3.7 80.4 $\pm$ 3.0 82.3 $\pm$ 3.8 81.7 $\pm$ 1.7 79.6 $\pm$ 0.7 70.1 $\pm$ 4.0
@@ 5.2. Results of Fairness Guidance
Generalization to Various GNN Architectures. In addition, to further verify the structural fairness of graphs guided by FairGuide, we evaluate the impact of fairness-guided links on two different GNN backbones, i.e. , GraphSage (hamilton2017inductive, ) and APPNP (gasteiger_predict_2019, ). Results in fairness are presented in Fig. 3. From Fig 3, we observe that different GNN backbones consistently achieve fairness improvements when trained on graphs guided by FairGuide. This demonstrates that FairGuide can grow fair graph structures that facilitate different GNN models.
@@ 5.5. Hyperparameter Analysis
In FairGuide, the cross-group boost rate $\beta$ and the number of community clusters C are two key hyperparameters.To investigate how the two hyperparameters affect fairness and task performance, We conduct GCN node classification task on the Pokec-n. Similar trends are also observed in other datasets. We vary $\beta$ to {0.01,0.1,1.0,10.0,100.0} and $C$ to {5,10,15,20,25}, respectively. All other settings follow the description given in Sec. 5.1. As shown in Fig. 7, the task performance F1 score remains stable across the tested hyperparameter ranges, fluctuating by only around 1–2 percentage points. For the fairness performance, FairGuide achieves a strong debiasing effect when $C\geq 15$ and $\beta\geq 1.0$. A larger number of clusters $C$ allows the model to capture more fine-grained feature information, and a large cross-group boost rate $\beta$ effectively establishes connections between different sensitive groups, which will benefit the graph structural fairness.